# ChatCFD: An LLM-Driven Agent for End-to-End CFD Automation with Domain-Specific Structured Reasoning

### v2

@@ Abstract
Computational Fluid Dynamics (CFD) is essential for advancing scientific and engineering fields, but is hindered by operational complexity, high expertise requirements, and limited accessibility. This paper introduces ChatCFD, an automated agent system for OpenFOAM simulations that processes multi-modal inputs (e.g., research papers, meshes) via an interactive interface, leveraging DeepSeek-R1 and DeepSeek-V3 large language models (LLMs), a multi-agent architecture, and OpenFOAM knowledge. Its four-stage pipeline, Knowledge Base Construction, User Input Processing, Case File Generation, and Execution and Error Reflection, enables iterative trial-reflection-refinement for intricate setups, supporting diverse physical models and external meshes. Validation on 205 benchmark tutorial cases, 110 perturbed variants, and 2 literature-derived cases demonstrates ChatCFD’s 82.1% operational success rate on basic cases, significantly outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), and 60–80% on literature-derived complex cases, with turbulence model studies showing a 40% success rate for common models vs. 10% for rare ones like RNGk-$\epsilon$. Physics coupling analyses reveal higher resource demands for multi-physics-coupled cases, while LLM bias toward simpler setups introduces persistent errors, such as dimensional inconsistency. Ablation studies highlight the efficacy of RAG-based modules and reflection mechanisms. By automating hypothesis testing and parameter exploration, ChatCFD accelerates scientific discovery in fluid mechanics and engineering, addressing LLM limitations through structured design and demonstrating strong potential as a modular component in MCP-based agent networks for collaborative multi-agent systems, paving the way for scalable AI-driven CFD innovation. The code for ChatCFD is available at: https://github.com/ConMoo/ChatCFD.
@@ 1 Introduction
Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [1, 2], energy systems [3, 4], urban environment [5, 6, 7], combustion [8, 9], and biomedical applications [10, 11]. It provides indispensable tools for simulating intricate fluid behaviors, thereby enabling design innovation and scientific discovery [12, 13]. However, CFD implementation faces substantial barriers, requiring deep domain expertise [14] and often relying on expensive commercial software. Even experts invest significant time in tasks such as solver selection, model setup, mesh generation, boundary condition definition, and post-processing [15]. These demands, combined with high computational costs, limit CFD accessibility for smaller organizations and hinder broader innovation [13]. Thus, there is an urgent need for automated, intuitive, and affordable CFD solutions. To meet this need, we present ChatCFD, an AI-driven agent system that streamlines CFD workflows, facilitating automated scientific discovery in fluid mechanics and engineering by enabling rapid iteration and exploration of complex flow phenomena without extensive manual intervention.
@@ 2.1 Stage 0: Knowledge Base Construction
The preprocessing aims to establish a foundational knowledge base for CFD tasks, performed in advance to optimize ChatCFD’s operations. Ideally, this knowledge base contains general CFD principles (e.g., numerical differentiation, discretization schemes), OpenFOAM-specific manuals, and a comprehensive example library. Fine-tuning large language models (LLMs) for general CFD knowledge is beyond the current scope; readers are referred to recent work by Dong et al. [30]. This study focuses on preprocessing OpenFOAM manuals and publicly accessible examples. A systematic analysis of OpenFOAM manuals and tutorials constructs a structured JSON database, cataloging essential parameters such as solver configurations, turbulence models, and file dependencies, providing a critical foundation for accurate case setup and error correction.
@@ 2.4 Stage 3: Error Correction and Reflection
  • ReferenceRetriever: This module retrieves reference files from the preprocessed OpenFOAM knowledge base, matching solver and model specifications from benchmark tutorial cases. If no exact match is found, it prioritizes solver-compatible files. To balance guidance and prompt complexity, two reference files are selected to aid error detection and correction, leveraging the LLM’s few-shot learning to enhance ChatCFD’s accuracy in error localization and resolution.
@@
  • Error Localization Module: Comprising two components: (1) DeepSeek-V3 rapidly identifies suspicious files based on error messages, reducing the analysis burden; (2) DeepSeek-R1 performs detailed reasoning to pinpoint erroneous files by comparing their content against the case library, leveraging its advanced inference capabilities for precise localization.
@@
  • Reflection Module: Activated during persistent errors, this module collects error messages and file modification histories, formulating reflections in the format: ”For situation A, I considered B but overlooked C; next, I will apply D to resolve the issue.” These reflections are stored as reflection blocks within a <reflexion> tag in the reflection history, enhancing iterative error correction by integrating with localization and correction modules.
@@ 3 Results and discussion
3.1 Performance Evaluation on OpenFOAM Basic Cases
@@
3.2 Evaluation and Analysis on Literature-Derived CFD Cases
@@
3.2.1 Ablation Study on ChatCFD System Configurations
@@
For the Nozzle Case 5, Figure 6 demonstrates a significant enhancement in the 10-step operational success rate with the integration of the ContextRetriever module. Configurations A through C, lacking ContextRetriever, achieved a 0% success rate for this metric, defined as completing 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors. Configuration D, incorporating ContextRetriever, improved this rate to 60%. This advancement stems from ContextRetriever’s ability to diagnose complex coupling-related errors within the MCP framework, addressing a key limitation of earlier configurations. For example, errors in thermophysical parameters, such as compressibility ($\Psi$), were often misattributed to the constant/thermophysicalProperties file in prior configurations. In reality, these errors frequently arise from inconsistencies in equation of state specifications within boundary condition entries in field files (e.g., 0/p, 0/T). By analyzing all relevant directories (0/, constant/, system/), ContextRetriever enables precise error localization and effective correction, enhancing performance for complex cases.
@@
3.2.2 Impact of Turbulence Models on ChatCFD Performance
@@ 4 Conclusion
This paper introduces ChatCFD, an innovative LLM-based CFD agent system for automating OpenFOAM simulations. By incorporating a multi-agent architecture, domain-specific knowledge bases, and advanced error reflection mechanisms, ChatCFD achieves end-to-end automation for reproducing literature results, including both incompressible (NACA0012 airfoil) and compressible (supersonic nozzle) flows. The simulations accurately reproduce velocity contours, pressure coefficients, and schlieren images, aligning closely with experimental data. ChatCFD’s innovations include: (1) an interactive multi-modal input mechanism integrating PDFs, meshes, and dialogue with physical constraints, providing a ChatGPT-like interface for guided literature analysis, case selection, and parameter specification, reducing operational barriers; (2) specialized AI agents leveraging CFD expertise for task interpretation, feature library construction via OpenFOAM preprocessing, and iterative reflection-refinement using DeepSeek-R1 for reasoning and DeepSeek-V3 for precise modifications; and (3) a robust error reflection system with ReferenceRetriever (tutorial templates) and ContextRetriever (case configurations) for real-time error classification and resolution.
Experimental validations on 315 basic cases (205 benchmark tutorials, 110 perturbed variants) yield an 82.1% operational success rate, outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), with lower success in multiphase/combustion due to interdisciplinary complexity. On literature-derived cases, ChatCFD achieves 60–80% success, with ablation studies showing comprehensive interpretation boosts accurate configuration from 0% to 40% for NACA0012 and 30% for Nozzle, while ContextRetriever improves operational success from 0% to 60% in Nozzle by resolving coupled-physics errors. Turbulence model analyses reveal 40% success for common models (e.g., Spalart-Allmaras) vs. 10% for rare ones (e.g., RNG k-$\epsilon$), attributed to data scarcity. Physics coupling studies indicate compressible flows demand tripled resources due to interdependencies, and LLM bias toward incompressible setups causes dimensional errors in complex cases. These findings highlight ChatCFD’s robustness, efficiency, and cost-effectiveness, establishing benchmarks for AI-driven CFD.
@@
ChatCFD effectively addresses the critical gaps in CFD automation agents. First, it integrates domain-specific knowledge through structured RAG and expert-designed agents, mitigating LLM training scarcity to enable precise boundary and initial condition specifications. Second, its modular, MCP-compatible design supports collaborative multi-agent networks, integrating specialized meshing agents to handle complex geometries beyond blockMesh. Third, by leveraging extensive literature-derived corpora with detailed specifications and benchmarks, it will facilitate scalable testing and iterative agent evolution, harnessing millions of CFD papers for refinement and validation in the future. These advancements accelerate scientific discovery in fluid mechanics and engineering by automating hypothesis testing, parameter optimization, and multiphysics simulations.

### v3

# ChatCFD: An LLM-Driven Agent for End-to-End CFD Automation with Structured Knowledge and Reasoning

@@ Abstract
Computational Fluid Dynamics (CFD) is essential for advancing scientific and engineering fields, but is hindered by operational complexity, high expertise requirements, and limited accessibility. This paper introduces ChatCFD, a LLM-driven agent system for end-to-end CFD automation. Powered by DeepSeek-R1/V3, a multi-agent architecture, structured OpenFOAM knowledge bases, precise error locator, and iterative reflection, ChatCFD dramatically outperforms prior systems. On 315 benchmark cases it attains 82.1% execution success (vs. 6.2% MetaOpenFOAM, 42.3% Foam-Agent) and, crucially, 68.12% physical fidelity—the first rigorous metric capturing whether a runnable simulation is scientifically meaningful. A dedicated Physics Interpreter achieves 97.4% summary fidelity, exposing the striking LLM gap between fluent narration and enforcing dozens of tightly coupled physical constraints in executable code. Resource-efficiency analysis highlights ChatCFD’s practical advantage, achieving an average of 192.1k tokens and \$0.208 per case—half the tokens of Foam-Agent and 1.5× cheaper than MetaOpenFOAM. Ablation studies confirm structured domain knowledge and reasoning is indispensable: removing the Solver Template DB collapses accuracy to 48%, while the Error Locator Module proves the single most critical component. Flexibility experiments show autonomous solver selection across compressible/incompressible and steady/transient regimes (95.23% success) and turbulence-model switching (100% success), even on unseen configurations. By faithfully reproducing complex literature cases NACA0012 airfoil and supersonic nozzle with 60–80% end-to-end success where baselines fail entirely, ChatCFD establishes new benchmarks for AI-driven CFD automation. Physics coupling analyses reveal higher resource demands for multi-physics-coupled cases, while LLM bias toward simpler setups introduces persistent errors. ChatCFD’s modular, MCP-compatible design directly enables collaborative multi-agent networks and paves the way for scalable AI-driven CFD innovation. The code for ChatCFD is available at: https://github.com/ConMoo/ChatCFD.
@@ 1 Introduction
Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [28, 10], energy systems [15, 2, 36], urban environment [20, 17, 26], combustion [24, 39, 33], astrophysics[40], and biomedical applications [22, 9]. It provides indispensable tools for simulating intricate fluid behaviors, thereby enabling design innovation and scientific discovery [3, 4]. However, CFD implementation faces substantial barriers, requiring deep domain expertise [32] and often relying on expensive commercial software. Even experts invest significant time in tasks such as solver selection, model setup, mesh generation, boundary condition definition, and post-processing [18]. These demands, combined with high computational costs, limit CFD accessibility for smaller organizations and hinder broader innovation [4]. Thus, there is an urgent need for automated, intuitive, and affordable CFD solutions. To meet this need, we present ChatCFD, an AI-driven agent system that streamlines CFD workflows, facilitating automated scientific discovery in fluid mechanics and engineering by enabling rapid iteration and exploration of complex flow phenomena without extensive manual intervention.
@@ 2 ChatCFD Pipeline Design
Figure 3: Based on the knowledge base constructed from the contents such as the OpenFOAM manual and OpenFOAM tutorials, and how it functions within the ChatCFD framework
@@ 2.1 Stage 0: Knowledge Base Construction
The preprocessing aims to establish a foundational knowledge base for CFD tasks, performed in advance to optimize ChatCFD’s operations. Ideally, this knowledge base contains general CFD principles (e.g., numerical differentiation, discretization schemes), OpenFOAM-specific manuals, and a comprehensive example library. Fine-tuning large language models (LLMs) for general CFD knowledge is beyond the current scope; readers are referred to recent work by Dong et al. [8]. This study focuses on preprocessing OpenFOAM manuals and publicly accessible examples.
Through a systematic analysis of OpenFOAM manuals and tutorials, we construct a structured JSON database that records key parameters such as solver configurations, turbulence models, and file dependencies, laying the foundation for accurate case setup and error correction. Additionally, a vector database is established to further enhance the understanding of OpenFOAM and to generate more precise OpenFOAM case descriptions. As shown in Fig. 3, the database primarily comprises two components: a Structured Database in JSON format and a Vector Database for efficient retrieval. The Structured Database includes the following parts:
  • File Dependency and Structure DB (db1): Designed to construct a hierarchical file structure model. This model is based on OpenFOAM tutorial cases and categorized primarily by solver type and turbulence model to store the necessary file structure for specific combinations. It also records the actual file structure of each tutorial case, which is utilized to construct and validate the space of permissible configurations.
  • Boundary Condition DB (db2): Comprehensively catalogs all available boundary condition types within OpenFOAM. For each boundary condition, it precisely documents the required additional setting keywords and their configuration specifications, ensuring the completeness and validity of the boundary setup.
  • Parameter Dimension DB(db3): Systematically records the standard dimensional configuration rules for all physical quantities in OpenFOAM. This database further defines the dimensional requirements and consistency constraints for relevant field variables under varying physical problems or specific solving environments.
  • Solver Template DB (db4): Systematically stores the core file content and structured metadata of fundamental OpenFOAM tutorial cases. The metadata includes: the case’s designated solver, turbulence model, and physical problem category. This information serves as the key basis for achieving high-accuracy classification and rapid screening of tutorial cases, thus providing precise reference templates for automated case initialization and configuration.
All configuration files are carefully parsed, converted to JSON format, and tagged with metadata (e.g., solver specifications and flow characteristics) for easy retrieval and application. OpenFOAM-specific header files are removed, and cases with external dependencies or auxiliary folders are excluded to maintain case autonomy and reduce the influence of irrelevant configuration elements on the LLM’s response quality. Furthermore, for cases where non-uniform field definitions occupy a large volume of file content (potentially exceeding dialogue context limits), placeholders (e.g., __CELL_COUNT__, __FACE_COUNT__, __SCALAR_DATA_PLACEHOLDER__) are used to replace the non-uniform field data. This allows the LLM to focus on critical settings (such as boundary and initial conditions) while mitigating the issue of a reduced case count due to complex data.
@@ 2.4 Stage 3: Error Correction and Reflection
  • ReferenceRetriever: This module retrieves reference files from the preprocessed OpenFOAM knowledge base, matching solver and model specifications from benchmark tutorial cases. If no exact match is found, it prioritizes solver-compatible files. To balance guidance and prompt complexity, two reference files are selected to aid error detection and correction, leveraging the LLM’s few-shot learning to enhance ChatCFD’s accuracy in error locator and resolution.
@@
  • Error Locator Module: Comprising two components: (1) DeepSeek-V3 rapidly identifies suspicious files based on error messages, reducing the analysis burden; (2) DeepSeek-R1 performs detailed reasoning to pinpoint erroneous files by comparing their content against the case library, leveraging its advanced inference capabilities for precise error location.
@@
  • Reflection Module: Activated during persistent errors, this module collects error messages and file modification histories, formulating reflections in the format: ”For situation A, I considered B but overlooked C; next, I will apply D to resolve the issue.” These reflections are stored as reflection blocks within a <reflexion> tag in the reflection history, enhancing iterative error correction by integrating with error locator and correction modules.
@@
2.5 Stage 4: Post-Processing
To enhance the transparency and user intelligibility of the computation process, ChatCFD integrates a dedicated Physics Interpreter in the fourth stage of the workflow. This component, implemented using the DeepSeek-R1 model, automatically produces a technical summary immediately after the case is successfully generated and deemed executable. The output is presented in Markdown format, supplemented by structured expressions such as tables and file tree diagrams. Its content systematically covers multiple dimensions, including fundamental information, physical problem background, special configuration settings, file structure, and key file contents, thereby comprehensively ensuring the transparency of the simulated physical process for the user.
@@ 3 Results and discussion
3.1 Benchmarking Execution Success Rates Across Agents
@@
3.2 Physical Fidelity Evaluation and Semantic Error Analysis
In Section 3.1, we assessed ChatCFD, MetaOpenFOAM, and Foam-Agent using success rate as the primary metric. While this allows rapid quantitative evaluation, successful execution does not ensure physical meaningfulness in CFD. To address this, we introduce physical fidelity as a core metric, evaluating whether executed simulations accurately capture user-specified physics. We also re-examine the link between run success and physical fidelity via agent-generated cases.
We define physical fidelity through a three-tier protocol. A case qualifies only if it meets all criteria without simplifications that compromise the scientific objective or user requirements:
  • Conditions Verification: Initial and boundary conditions match the user-described scenario.
  • Model Confirmation: Physical properties, models, and parameters are appropriate and consistent.
  • Flow Feature Comparison: Post-processed visualizations align with ground-truth flow characteristics.
Using this protocol, we analyzed ChatCFD and Foam-Agent results (Fig. 6). ChatCFD’s physical fidelity rate among runnable cases (phy|run) is lower than Foam-Agent’s, but its overall rate is higher, confirming superior performance. Note that phy|run, limited to runnable cases, may inflate for agents handling only simple setups.
Figure 6: Physical fidelity evaluation of ChatCFD. (a) Success rates of ChatCFD and Foam-Agent under three criteria: run — execution success (simulation completes without crashing); $\midrun — phy$sical fidelity conditioned on execution success (proportion of runnable cases that are physically correct); phy — overall physical fidelity (proportion of all cases that are both runnable and physically meaningful). (b) Root-cause analysis of semantic failures in ChatCFD-generated cases that execute successfully but lack physical fidelity, categorized by error type (boundary/initial conditions, physical properties/models, numerical schemes, etc.).
Additionally, we summarized the causes of physical incorrectness in runnable cases. The results, shown in Fig 6, indicate that incorrect settings of boundary conditions, initial conditions, and initial fields in multiphase flows are the main contributors, accounting for 60 % of the issues, as these require deep physical insight and geometric awareness. Such misconfigurations rarely trigger convergence failures, allowing apparent success that misleads agents. Similarly, omissions in setFieldsDict often evade detection, exacerbating infidelity.
To further enhance transparency, we integrated a Physics Interpreter, evaluated using DeepSeek-V3 and GPT-4o for summary rationality. It achieved 97.4% fidelity, confirming accurate natural-language conveyance of the agent’s intended physics.
The contrast between the near-perfect fidelity of the Physical Interpreter (97.4%) and the overall Physical Fidelity of only 68.12% is particularly interesting. This suggests that while a seemingly simple user request (e.g., "2-D flow around a cylinder at Re = 100") can be almost perfectly restated in natural language by the LLM, its faithful implementation in executable OpenFOAM code requires dozens of tightly interdependent, domain-specific settings (such as inlet velocity profile, turbulence suppression, kinematic viscosity, blockage ratio, spanwise boundary conditions, etc.). This highlights a key limitation of current LLMs: they excel at high-level linguistic narration but consistently struggle to enforce the complete chain of physical constraints within the generated executable code.
3.3 Ablation Study on ChatCFD Architectural Components
To quantitatively assess the contribution of the proposed database system and functional modules to the overall system performance, we conducted an ablation study using $20\%$ of the cases sampled from the OpenFOAM benchmark test set. It should be noted that the degree of influence exerted by different components will vary due to the inherent differences in task types. Consequently, for the specific task of literature replication, we will perform additional ablation experiments in Section 3.5.1 for further, dedicated analysis.
The first four experiments were designed to evaluate the impact of the knowledge databases on ChatCFD’s performance. This knowledge base primarily influences Stage2 and Case File Generation and Error Correction and Reflection in Stage3. Simultaneously, the experimental groups "w/o mod1" (removal of the Error Locator Module) and "w/o mod2" (removal of the Reflection Module) were used to verify and quantify the efficacy of the error locator and reflection mechanisms, with their influence predominantly focused on Error Correction and Reflection in Stage3.
The impact of the first four knowledge base ablation experiments on ChatCFD’s performance, ranked from highest to lowest influence, is as follows: Solver Template DB (db4, 48 %), Boundary Condition DB (db2, 57 %), Parameter Dimension DB (db3, 68 %), and File Dependency and Structure DB (db1, 73 %). The results show that the Solver Template DB is utilized across all operational stages of ChatCFD and is crucial for leveraging the Large Language Model’s few-shot learning capability to generate initially correct case files.
Figure 7: Ablation study quantifying the contribution of individual prior-knowledge components and core modules to ChatCFD’s configuration accuracy. - w/o File Dependency DB (DB1): removes the structured JSON database that enforces mandatory file hierarchy and dictionary dependencies. - w/o Boundary Condition DB (DB2): disables the boundary-condition template database (no validated keyword requirements). - w/o Parameter Dimension DB (DB3): disables automatic dimension enforcement and file-format checks. - w/o Solver Template DB (DB4): removes all reference tutorial snippets and metadata. - w/o Error Locator Module (Module 1): disables the error-positioning module. - w/o Reflection (Module 2): disables the iterative reflection loop. - w/ all: full ChatCFD system (all four prior-knowledge databases + all function modules).
Fig 7 shows that the removal of the Error Locator Module (w/o mod1) resulted in the lowest success rate. Here, we bypassed it and prompted DeepSeek-R1 to modify files based solely on error messages. The experimental results reveal two findings: 1) Accurate identifying of the erroneous file is a critical step for enhancing the agent’s efficiency during the reflection and correction phase (Stage 3); and 2) Relying only on error messages is ineffective. Its shortcomings are particularly pronounced when dealing with errors that do not explicitly point to the source file, such as dimensional errors. Removing the Reflection Module (w/o mod2) causes the smallest performance drop because it primarily benefits high-difficulty cases, which constitute only a minor fraction of the benchmark. In contrast, the other components—knowledge bases and Error Locator—impact cases across all difficulty levels.
3.4 Flexibility Evaluation of Alternative OpenFOAM Configurations
To evaluate ChatCFD’s ability to adapt to different physical assumptions and numerical settings, we designed two targeted experiments:
  • Solver selection across regimes — Given only a textual description, ChatCFD autonomously determines the flow regime (compressible/incompressible, steady/transient) and selects the appropriate solver.
  • Turbulence-model switching — With mesh and boundary conditions fixed, only the turbulence closure is changed, requiring automatic rewriting of constant/turbulenceProperties and related dictionaries.
Fig 8 summarizes the results of Experiment 1. ChatCFD correctly identified the required solver in 20 of 21 cases, including the unseen “simpleCar” compressible-transient configuration (not present in the database). The sole failure occurred on “buoyantCavity” under compressible-transient assumptions, which currently exceeds ChatCFD’s supported scope. In the turbulence-switching experiment (Figure 9), ChatCFD successfully applied $\epsilon, k$-$\omega$ SST, and laminar closures to every case—including the originally laminar damBreak—demonstrating robust handling of non-default configurations. These results confirm ChatCFD’s strong generalization: it is not limited to tutorial-default settings but can reason about physically valid alternatives, significantly enhancing its practical utility for practical CFD exploration.
Figure 8: Solver flexibility under varying physical regimes. Blue circles indicate the reference (tutorial) setting; green checks mark successful execution with the identified solver; red crosses denote failure. ChatCFD correctly selects the appropriate solver in 20/21 of cases, including unseen compressible-transient scenarios.
Figure 9: Turbulence-model flexibility. ChatCFD successfully switches between $\epsilon, k$-$\omega$ SST, and other closures for all tested cases, including the originally laminar damBreak (100% success). Symbols follow Figure 8.
3.5 Evaluation of CFD Cases Reproduced from Literature
@@
3.5.1 Ablation Study on ChatCFD Configurations for Literature Reproduction
@@
For the Nozzle Case 5, Figure 11 demonstrates a significant enhancement in the 10-step operational success rate with the integration of the ContextRetriever module. Configurations A through C, lacking ContextRetriever, achieved a 0% success rate for this metric, defined as completing 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors. Configuration D, incorporating ContextRetriever, improved this rate to 60%. This advancement stems from ContextRetriever’s ability to diagnose complex coupling-related errors within the MCP framework, addressing a key limitation of earlier configurations. For example, errors in thermophysical parameters, such as compressibility ($\Psi$), were often misattributed to the constant/thermophysicalProperties file in prior configurations. In reality, these errors frequently arise from inconsistencies in equation of state specifications within boundary condition entries in field files (e.g., 0/p, 0/T). By analyzing all relevant directories (0/, constant/, system/), ContextRetriever enables precise error locator and effective correction, enhancing performance for complex cases.
@@
3.5.2 Influence of Turbulence Models on ChatCFD Efficacy
@@ 4 Conclusion
This paper introduces ChatCFD, an innovative LLM-based CFD agent system for automating OpenFOAM simulations. By integrating a multi-agent architecture, domain-specific structured knowledge bases, precise error locator, and iterative reflection, ChatCFD dramatically outperforms prior systems while exposing fundamental limitations of current LLMs in scientific computing.
On 315 OpenFOAM benchmark cases, ChatCFD attains an 82.1% execution success rate—far surpassing MetaOpenFOAM (6.2%) and Foam-Agent (42.3%). More importantly, we introduced physical fidelity as a rigorous new metric: among all attempts, ChatCFD generates physically correct solutions in 68.12% of cases, with 60% of semantic failures traced to subtle boundary/initial condition errors that escape convergence checks. A dedicated Physics Interpreter further achieves 97.4% summary fidelity, revealing the striking gap between LLMs’ fluency in high-level narration and their struggle to enforce dozens of tightly coupled domain constraints in executable code.
Resource-efficiency analysis on 205 benchmark cases further underscores ChatCFD’s practical superiority. It consumes only 192.1k tokens and \$0.208 per case—roughly half the tokens of Foam-Agent and 1.5× cheaper than MetaOpenFOAM—while requiring the fewest reflection iterations on average. This efficiency gain over naive LLM prompting stems from strategic model routing (DeepSeek-V3 for generation, DeepSeek-R1 for reasoning) and a reflection memory that filters irrelevant context. Even under a strict iteration limit, ChatCFD’s reflection trajectory consistently outperforms baselines, solving cases that saturate the others at  50% success after 25 iterations. These results demonstrate that structured domain integration not only boosts accuracy but dramatically reduces the operational cost of LLM-driven scientific computing, making high-fidelity CFD automation viable at scale.
Ablation studies confirm that structured OpenFOAM knowledge is indispensable: removing the Solver Template DB alone collapses accuracy to 48%, while the Error Locator Module proves the single most critical component for real-world robustness. Flexibility experiments demonstrate ChatCFD’s ability to autonomously select appropriate solvers across compressible/incompressible and steady/transient regimes (95.23% success) and switch turbulence closures (100% success), even on unseen configurations.
By reproducing complex literature cases—NACA0012 airfoil, supersonic nozzles—ChatCFD shows 60–80% end-to-end success where baselines fail entirely. These results establish potentially new benchmarks for AI-driven CFD and provide the quantitatively grounded evidence of where current LLMs break in scientific tool use.
@@
ChatCFD effectively addresses the critical gaps in CFD automation agents. First, it integrates domain-specific knowledge through structured RAG and expert-designed agents, mitigating LLM training scarcity to enable precise boundary and initial condition specifications. Second, its modular, MCP-compatible design supports collaborative multi-agent networks, integrating specialized meshing agents to handle complex geometries beyond blockMesh. Third, by leveraging extensive literature-derived corpora with detailed specifications and benchmarks, it will facilitate scalable testing and iterative agent evolution, harnessing millions of CFD papers for refinement and validation in the future. These advances will enable scientists across disciplines to explore bold ideas using CFD that were previously stalled by technical details. By liberating researchers from low-level case-setup burdens, it enables non-experts to perform computational fluid dynamics studies, allows rapid exploration of hypothesis and design spaces, and frees CFD specialists to focus on genuine scientific questions.